OpenStudy (anonymous):

what is the general form of an equation with a point at (-10,4) and that has an undefined slope???

OpenStudy (mokeira):

if slope is undefined, m=0 \[m=\frac{ \Delta y }{ \Delta x }\] \[0=\frac{ y-4 }{x--10 }\] 0=y-4 y=4 that is the equation @kspik